OpenStudy (anonymous):

Find the exact value of sin157.5 in half angle identities

OpenStudy (anonymous):

kk ! your answer :) ?!

OpenStudy (anonymous):

i will send it to you right now

OpenStudy (anonymous):

kk !

OpenStudy (anonymous):

i need to find it in half angle identities

OpenStudy (anonymous):

you're right ! @lsugano

OpenStudy (anonymous):

but the answer book, it said \[\sqrt{\left(\begin{matrix}2-\sqrt{2} \\ 2\end{matrix}\right)}\]

OpenStudy (anonymous):

i cant get this answer

OpenStudy (anonymous):

is this what you meant : \[\sqrt{\frac{2-\sqrt{2}}{4}}\]

OpenStudy (anonymous):

yes! sorry i dont know how to use this equation bar haha

OpenStudy (anonymous):

the answer is that but i cant get it

OpenStudy (anonymous):

in your attached pic you got it .....well almost. just divide :D

OpenStudy (anonymous):

like this ?

OpenStudy (anonymous):

Exactly!

OpenStudy (anonymous):